# Restrictions on Initial Conditions in Cosmological Scenarios and Implications for Simulations of Primordial Black Holes and Inflation

### v1

@@ I Introduction
Specific examples of cosmological processes that cannot be studied within the framework of perturbation theory include the possible formation of primordial black holes [1, 2, 3, 4] (PBHs) and the robustness of inflationary models to inhomogeneities in the early universe. Understanding whether, how, and where PBHs form [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] provides insights into the fundamental physics of inflation [35, 36, 37, 38], the nature of dark matter [39, 40, 41, 42, 43], and the possible origin of supermassive black holes [44, 45]. Similarly, whether or not inflation, in the presence of severe early-time inhomogeneity, remains an attractor [46, 47] has also been investigated non-linearly [48, 49, 50, 51, 52, 53, 54, 55, 56], with one thorny question being what form of “generic” initial data is the right one to test such a scenario. Since most of our cosmological intuition is anchored to the FLRW background, there is a tendency to imagine the non-linear physics exists on top of the underlying cosmology. However, the non-linear nature of Einstein’s equations does not generally allow for such a separation. In order to study these and other dynamic processes in cosmology, we have to solve the equations—including those for any initial conditions—self-consistently, and this may result in a solution very far from the FLRW one onto which any perturbations were added.
@@
Specifically, we apply a simple but analytical toy model, previously discussed in [62], to a cosmological context in order to demonstrate surprising consequences of the non-linearities in the Hamiltonian constraint. In particular, we show that solutions to the Hamiltonian constraint exist only for certain combinations of the deviations of the mass-energy density and extrinsic curvature from the asymptotic cosmology. For example, assuming that the extrinsic curvature remains unchanged from the cosmological background will allow for density inhomogeneities only up to a surprisingly small maximum value for horizon scale modes. Stated differently, this implies that the intrinsic curvature alone cannot balance arbitrarily large inhomogeneities in the mass-energy density. Moreover, we demonstrate that, if solutions exist, they are not unique. Instead, there exist two branches of solutions for allowed combinations of the deviations in the mass-energy density and extrinsic curvature. One of these branches corresponds to solutions with weak intrinsic curvature and approaches the FLRW spacetime in the limit of small inhomogeneities, while the second branch corresponds to strong intrinsic curvature and may feature black-hole horizons and throat-like structures in the initial data. While both branches represent viable initial data in the sense that they satisfy the Hamiltonian constraint, they may or may not not be astrophysically realistic, in the sense of describing a time-like slice of the universe as it may have emerged from its prior evolution.
@@ II A toy model for the Hamiltonian constraint
In the context of a 3+1 decomposition we identify hypersurfaces of constant coordinate time $t$ with a spatial slice. We denote the normal vector on these slices with $n^{a}$ and the induced spatial metric with $\gamma_{ab}=g_{ab}+n_{a}n_{b}$, where $g_{ab}$ is the spacetime metric. In the preferred coordinate system of an FLRW spacetime discussed above, the normal vectors $n^{a}$ coincide with the four-velocities $u^{a}$ of coordinate observers (see, e.g., Excercise 4.1 in [58]).
In the following we will focus in spherically symmetric spaces, which are always conformally flat, so that we may write $\gamma_{ij}=\psi^{4}\eta_{ij}$. Here $\psi$ is the conformal factor and $\eta_{ij}$ the flat metric. For the preferred slicing of a FLRW spacetime we may then identify $\psi=a^{1/2}$ (with $\psi=1$ on the initial slice at time $t_{0}$); moreover, the mean curvature $K$, i.e. the trace of the extrinsic curvature $K=\gamma^{ij}K_{ij}$, which is related to the time-derivative of the spatial metric, is given by
$K_{0}=-3H=-3\dot{a}/a=(24\pi\rho_{0})^{1/2}.$ (2)
We now set up a scenario in which a local, spherically symmetric over-dense region is embedded at the center of an otherwise homogeneous cosmological spacetime. Specifically, we consider an initial (spatial) slice at coordinate time $t=0$, say, also assumed to be spherically symmetric. At large distances from the center we choose the slice to approach that of an FLRW spacetime, i.e. with $K\rightarrow(24\pi\rho_{0})^{1/2}$ and $\psi\rightarrow 1$. Following [62] we explore properties of solutions to the Hamiltonian constraint by adopting a toy model with a constant over-density inside some (coordinate) radius $r_{\rm od}$; i.e. we assume that the mass-energy density $\rho=n_{a}n_{b}T^{ab}$ is given by
@@
Here $\rho_{0}$ is the density of the corresponding asymptotic FLRW spacetime, and $\Delta\rho$ the over-density. We make no assumptions on the size of $R$; the over-dense region may or may not be fully inside the Hubble volume, $\sim H^{-3}$, of the associated asymptotic spacetime. The Hamiltonian constraint then becomes
@@
where $\bar{\nabla}^{2}$ is the flat-space Laplace operator and $\bar{A}_{ij}$ the conformally related trace-free part of the extrinsic curvature.
@@
is no longer satisfied analytically with $\bar{A}_{ij}=0$ and a vanishing momentum density, $S^{i}=0$. In the following we assume that the contributions from the derivative of $K$ are fully balanced by an appropriate choice of $S^{r}$, treating it as a fluid for which the momentum density and the energy density are independent degrees of freedom. We may therefore continue to choose $\bar{A}_{ij}=0$, and only concern ourselves with the Hamiltonian constraint. Allowing a non-vanishing $\bar{A}_{ij}$ would affect our solutions quantitively, but we expect our conclusions to hold qualitatively even in that case.
@@
Specifically, if the over-dense region is of Hubble scale, $n\simeq 1$, the maximum value of $\delta rho$ that can be accommodated by the intrinsic curvature alone, i.e. with $\delta K=0$, is given by
@@
Figure 2: Top panel: Representative solutions $\psi(r)$ for a fixed value of the additional expansion $\delta K=0.1$, a Hubble size region $n=0.5$ and the two values of $\alpha$ from the strong and weak branches as marked on Fig. 1, namely $\alpha=0.7$ and $\alpha=19.8$. Both values give $f(\alpha)=0.05$ and so the same profile for $\rho$, but the profiles for $\psi$ differ. The vertical lines show the coordinate locations of the cosmological horizons for the two solutions. Bottom panel: The proper radius $R=\psi^{2}r$ as a function of the coordinate radius $r$ for the two solutions. Although the coordinate radii of the cosmological horizons are different in each solution, both horizons have the same proper radius $R(r_{\rm cos})=1$ and therefore the same entropy. They are, however, physically distinct, with those on the strong branch exhibiting a non-monotonic behavior in the proper radius $R$ as a function of coordinate radius $r$ in cases with $\alpha<1$.
@@ III.1 Misner-Sharp Mass
for the compaction of the over-dense region, where $R_{\rm od}=\psi^{2}r_{\rm od}$ is the proper radius of the over-dense region. Note that $M_{\rm MS}/R_{\rm od}\rightarrow\infty$ in the limit $\alpha\rightarrow 0$, indicating that small values of $\alpha$ correspond to strong-field solutions, while large values of $\alpha$ correspond to weak-field solutions. In Fig. 3 we show values of the compaction (III.1) versus the effective energy density, $\delta\rho-\delta K(2+\delta K)$ (from (14)) for families with different values of $n$.
@@
In the bottom panel of Fig. 2 we also show the proper radius $R=\psi^{2}r$ as a function of the coordinate radius $r$ for the two representative examples of the strong-field and weak-field solutions high-lighted in Fig. 1. In particular, notice that the proper radius monotonically increases with $r$ for the weak-field solution, but features a non-monotonic behavior for the strong-field solution. In [66], this behavior was used to identify “Type-II” initial data, which suggests that the initial data used in that work belong to what we refer to as strong-field branch data.
@@ III.2 Horizons
We order to identify horizons we compute the expansion of both outgoing and ingoing null geodesics, which yields [^2] Note that the subscripts “out” and “in” refer to the direction of null geodesics, and not to the interior or exterior of the over-dense region.
@@
respectively (see, e.g., [67]). We then identify trapped surfaces with regions for which $\Theta_{(\rm{out})}<0$, and the black-hole apparent horizon $r_{\rm BH,out}$ with the marginally outermost trapped surface, i.e. the largest value of $r$ for which $\Theta_{(\rm{out})}=0$. We similarly identify a cosmological apparent horizon $r_{\rm cos}$ with surfaces for which $\Theta_{(\rm{in})}=0$.
We will also evaluate the entropy of the cosmological horizons [68]. Whilst this is a dynamical case, not a stationary one, and so the entropy is not well-defined, in spherical symmetry recent arguments have suggested that it may be approximated by the proper area of the apparent horizon (see e.g. [69])
@@ III.2.1 Black hole horizons
Exact solutions to the cubic equation (34) can also be found using, for example, the recipe on page 228 of [70]. In particular, this involves computing expressions $\bar{Q}$ and $\bar{R}$ from the coefficients of the cubic, which, in our case, reduce to [^3] We use bars on $\bar{Q}$ and $\bar{R}$ in order to avoid confusion with the proper radius $R$.
@@ III.2.2 Cosmological horizons
Even though this (coordinate) radius approaches zero as $\alpha\rightarrow 0$, the horizon’s proper radius does not. To see this, we compute, for any value of $\alpha$,
@@
Eq. (III.2.2) shows several interesting properties of the proper radius of the cosmological horizon, and hence its proper area and entropy. We first observe that $f(\alpha)\rightarrow 0$ as $\alpha\rightarrow 0$, so that $R_{\rm cos}$ indeed remains finite in this limit, as we stated above. The same is true in the weak-field limit $\alpha\rightarrow\infty$, for which $f(\alpha)\rightarrow 0$ also. We therefore see that
@@
In fact, the cosmological horizon’s proper radius (III.2.2) depends on $\alpha$ only through the function $f(\alpha)$ defined in (14). As we discussed there, a given value of $f(\alpha)$, corresponding to a given value of the effective energy density $\delta\rho-\delta K(2+\delta K)$, permits two different solutions for $\alpha$, defining the weak-field and strong-field branches of solutions. However, since (III.2.2) depends on $f(\alpha)$ only, and not on $\alpha$ individually, we see that the cosmological horizon’s proper radius, for given values of $\delta\rho$ and $\delta K$, takes the same value on the weak-field and strong-field branches, as does its associated entropy.

### v2

@@ I Introduction
Specific examples of cosmological processes that cannot be studied within the framework of perturbation theory include the possible formation of primordial black holes [1, 2, 3, 4] (PBHs) and the robustness of inflationary models to inhomogeneities in the early universe. Understanding whether, how, and where PBHs form [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] provides insights into the fundamental physics of inflation [36, 37, 38, 39], the nature of dark matter [40, 41, 42, 43, 44], and the possible origin of supermassive black holes [45, 46]. Similarly, whether or not inflation, in the presence of severe early-time inhomogeneity, remains an attractor [47, 48] has also been investigated non-linearly [49, 50, 51, 52, 53, 54, 55, 56, 57], with one thorny question being what form of “generic” initial data is the right one to test such a scenario. Since most of our cosmological intuition is anchored to the FLRW background, there is a tendency to imagine the non-linear physics exists on top of the underlying cosmology. However, the non-linear nature of Einstein’s equations does not generally allow for such a separation. In order to study these and other dynamic processes in cosmology, we have to solve the equations—including those for any initial conditions—self-consistently, and this may result in a solution very far from the FLRW one onto which any perturbations were added.
@@
Specifically, we apply a simple but analytical toy model, previously discussed in [63] and motivated by [64], to a cosmological context in order to demonstrate surprising consequences of the non-linearities in the Hamiltonian constraint. In particular, we show that solutions to the Hamiltonian constraint exist only for certain combinations of the deviations of the mass-energy density and extrinsic curvature from the asymptotic cosmology. For example, assuming that the extrinsic curvature remains unchanged from the cosmological background will allow for density inhomogeneities only up to a surprisingly small maximum value for horizon scale modes. Stated differently, this implies that the intrinsic curvature alone cannot balance arbitrarily large inhomogeneities in the mass-energy density. Moreover, we demonstrate that, if solutions exist, they are not unique. Instead, there exist two branches of solutions for allowed combinations of the deviations in the mass-energy density and extrinsic curvature. One of these branches corresponds to solutions with weak intrinsic curvature and approaches the FLRW spacetime in the limit of small inhomogeneities, while the second branch corresponds to strong intrinsic curvature and may feature black-hole horizons and throat-like structures in the initial data. While both branches represent viable initial data in the sense that they satisfy the Hamiltonian constraint, they may or may not not be astrophysically realistic, in the sense of describing a time-like slice of the universe as it may have emerged from its prior evolution.
@@ II A toy model for the Hamiltonian constraint
In the context of a 3+1 decomposition we identify hypersurfaces of constant coordinate time $t$ with a spatial slice. We denote the normal vector on these slices with $n^{a}$ and the induced spatial metric with $\gamma_{ab}=g_{ab}+n_{a}n_{b}$, where $g_{ab}$ is the spacetime metric. In the preferred coordinate system of an FLRW spacetime discussed above, the normal vectors $n^{a}$ coincide with the four-velocities $u^{a}$ of coordinate observers (see, e.g., Exercise 4.1 in [59]).
In the following we will focus on spherically symmetric spaces, which are always conformally flat, so that we may write $\gamma_{ij}=\psi^{4}\eta_{ij}$. Here $\psi$ is the conformal factor and $\eta_{ij}$ the flat metric. While this choice simplifies our analysis, we expect our findings to hold qualitatively even in the absence of spherical symmetry or conformal flatness (see, e.g., [64] for a numerical example, albeit for asymptotically flat spacetimes, that assumes neither spherical symmetry nor conformal flatness). For the preferred slicing of a FLRW spacetime we may then identify $\psi=a^{1/2}$ (with $\psi=1$ on the initial slice at time $t_{0}$); moreover, the mean curvature $K$, i.e. the trace of the extrinsic curvature $K=\gamma^{ij}K_{ij}$, which is related to the time-derivative of the spatial metric, is given by
$K_{0}=-3H=-3\dot{a}/a=-(24\pi\rho_{0})^{1/2}.$ (2)
We now set up a scenario in which a local, spherically symmetric over-dense region is embedded at the center of an otherwise homogeneous cosmological spacetime. Specifically, we consider an initial (spatial) slice at coordinate time $t=0$, say, also assumed to be spherically symmetric. At large distances from the center we choose the slice to approach that of an FLRW spacetime, i.e. with $K\rightarrow-(24\pi\rho_{0})^{1/2}$ and $\psi\rightarrow 1$. Following [63] we explore properties of solutions to the Hamiltonian constraint by adopting a toy model with a constant over-density inside some (coordinate) radius $r_{\rm od}$; i.e. we assume that the mass-energy density $\rho=n_{a}n_{b}T^{ab}$ is given by
@@
Here $\rho_{0}$ is the density of the corresponding asymptotic FLRW spacetime, and $\Delta\rho$ the over-density. We make no assumptions on the size of $r_{od}$; the over-dense region may or may not be fully inside the Hubble volume, $\sim H^{-3}$, of the associated asymptotic spacetime. The Hamiltonian constraint then becomes
@@
where $\bar{\nabla}^{2}$ is the flat-space Laplace operator and
$\bar{A}_{ij}=\psi^{2}A_{ij}=\psi^{2}\left(K_{ij}-\frac{1}{3}\gamma_{ij}K\right)$ (5)
the conformally related trace-free part of the extrinsic curvature
@@
In many numerical relativity applications the wrong-sign issue is avoided by introducing a conformal rescaling $\bar{\rho}=\psi^{n}\rho$ of the density, with $n\geq 5$, designed to flip the sign of the exponent of $\psi$ on the right-hand side of the Hamiltonian constraint (4), and specifying $\bar{\rho}$ rather than $\rho$. In that case, a solution for $\psi$ exists for any value of $\bar{\rho}$. However, if one reconstructs $\rho$ from $\bar{\rho}$ one finds that the limits on the size of $\rho$ have not changed from those presented in [63]. This effect is most important when the size of the over-dense region is large. Here we choose to study solutions of the Hamiltonian constraint for given values of the physical density, rather than the rescaled density, since that is the quantity that can be directly computed from the contents of the Universe.
@@
is no longer satisfied analytically with $\bar{A}_{ij}=0$ and a vanishing momentum density, $S^{i}=0$. In the following we assume that the contributions from the derivative of $K$ are fully balanced by an appropriate choice of $S^{r}$, treating it as a fluid for which the momentum density and the energy density are independent degrees of freedom. [^1] Using a fluid treatment simplifies the process of simultaneously solving the Hamiltonian and momentum constraints; in instances where the energy density $\rho$ and the momentum density $S_{i}$ are both functions of a common object, say a scalar field, one has to solve these constraints self-consistently. The physical density $\rho$ associated with a scalar field will still satisfy the Hamiltonian constraint, however, and therefore will still be subject to the “wrong-sign” issue discussed above. Accordingly, there will be restrictions on such initial data similar to those discussed here even in that case. We may therefore continue to choose $\bar{A}_{ij}=0$, and only concern ourselves with the Hamiltonian constraint. Allowing a non-vanishing $\bar{A}_{ij}$ would affect our solutions quantitatively, but, because the “wrong-sign” issue discussed above persists for non-zero $\bar{A}_{ij}$, we expect our conclusions to hold qualitatively even in that case (see also [64] for a numerical example).
@@
Specifically, if the over-dense region is of Hubble scale, $n\simeq 1$, the maximum value of $\delta\rho$ that can be accommodated by the intrinsic curvature alone, i.e. with $\delta K=0$, is given by
@@
Figure 2: Top panel: Representative solutions $\psi(r)$ for a fixed value of the additional expansion $\delta K=0.1$, a Hubble size region $n=0.5$ and the two values of $\alpha$ from the strong and weak branches as marked on Fig. 1, namely $\alpha=0.7$ and $\alpha=19.8$. Both values give $f(\alpha)=0.05$ and so the same profile for $\rho$, but the profiles for $\psi$ differ. The vertical lines show the coordinate locations of the cosmological horizons for the two solutions. Bottom panel: The areal radius $R=\psi^{2}r$ as a function of the coordinate radius $r$ for the two solutions. Although the coordinate radii of the cosmological horizons are different in each solution, both horizons have the same areal radius $R(r_{\rm cos})=1$ and therefore the same entropy. They are, however, physically distinct, with those on the strong branch exhibiting a non-monotonic behavior in the areal radius $R$ as a function of coordinate radius $r$ in cases with $\alpha<1$.
@@ III.1 Misner-Sharp Mass
for the compaction of the over-dense region, where $R_{\rm od}=\psi^{2}r_{\rm od}$ is the areal radius of the over-dense region. Note that $M_{\rm MS}/R_{\rm od}\rightarrow\infty$ in the limit $\alpha\rightarrow 0$, indicating that small values of $\alpha$ correspond to strong-field solutions, while large values of $\alpha$ correspond to weak-field solutions. In Fig. 3 we show values of the compaction (III.1) versus the effective energy density, $\delta\rho-\delta K(2+\delta K)$ (from (15)) for families with different values of $n$.
@@
In the bottom panel of Fig. 2 we also show the areal radius $R=\psi^{2}r$ as a function of the coordinate radius $r$ for the two representative examples of the strong-field and weak-field solutions high-lighted in Fig. 1. In particular, notice that the areal radius monotonically increases with $r$ for the weak-field solution, but features a non-monotonic behavior for the strong-field solution. In [68], this behavior was used to identify “Type-II” initial data, which suggests that the initial data used in that work belong to what we refer to as strong-field branch data.
@@ III.2 Horizons
In order to identify horizons we compute the expansion of both outgoing and ingoing null geodesics, which yields [^3] Note that the subscripts “out” and “in” refer to the direction of null geodesics, and not to the interior or exterior of the over-dense region.
@@
respectively (see, e.g., [69]). We then identify trapped surfaces with regions where inward and outward null expansions are negative. The black-hole apparent horizon $r_{\rm BH,out}$ is the marginally outer trapped surface, the outermost surface where the outgoing null expansion is zero, i.e. the largest value of $r$ for which $\Theta_{(\rm{out})}=0$. We similarly identify a cosmological apparent horizon $r_{\rm cos}$ with surfaces for which the ingoing null expansion $\Theta_{(\rm{in})}=0$. (See also [70] for a discussion of different types of horizon).
We will also evaluate the entropy of the cosmological horizons [71]. Whilst this is a dynamical case, not a stationary one, and so the entropy is not well-defined, in spherical symmetry recent arguments have suggested that it may be approximated by the proper area of the apparent horizon (see most recently [72], and key works on dynamical horizons [73, 74, 75])
@@ III.2.1 Black hole horizons
Exact solutions to the cubic equation (35) can also be found using, for example, the recipe on page 228 of [76]. In particular, this involves computing expressions $\bar{Q}$ and $\bar{R}$ from the coefficients of the cubic, which, in our case, reduce to [^4] We use bars on $\bar{Q}$ and $\bar{R}$ in order to avoid confusion with the areal radius $R$.
@@ III.2.2 Cosmological horizons
Even though this (coordinate) radius approaches zero as $\alpha\rightarrow 0$, the horizon’s areal radius does not. To see this, we compute, for any value of $\alpha$,
@@
Eq. (III.2.2) shows several interesting properties of the areal radius of the cosmological horizon, and hence its proper area and entropy. We first observe that $f(\alpha)\rightarrow 0$ as $\alpha\rightarrow 0$, so that $R_{\rm cos}$ indeed remains finite in this limit, as we stated above. The same is true in the weak-field limit $\alpha\rightarrow\infty$, for which $f(\alpha)\rightarrow 0$ also. We therefore see that
@@
In fact, the cosmological horizon’s areal radius (III.2.2) depends on $\alpha$ only through the function $f(\alpha)$ defined in (15). As we discussed there, a given value of $f(\alpha)$, corresponding to a given value of the effective energy density $\delta\rho-\delta K(2+\delta K)$, permits two different solutions for $\alpha$, defining the weak-field and strong-field branches of solutions. However, since (III.2.2) depends on $f(\alpha)$ only, and not on $\alpha$ individually, we see that the cosmological horizon’s areal radius, for given values of $\delta\rho$ and $\delta K$, takes the same value on the weak-field and strong-field branches, as does its associated entropy.